OpenStudy (anonymous):

Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $111.00 for 3 days and 300 miles, while Mary was charged $207.00 for 5 days and 600 miles.What does Best Rentals charge per day and per mile?

OpenStudy (anonymous):

first set up 2 equations:

OpenStudy (anonymous):

how would those two equations look like ?

OpenStudy (anonymous):

300m/3d=111 and 600m/5d=207

OpenStudy (anonymous):

now what you do is isolate one variable on one equation so that you can substitute

OpenStudy (anonymous):

what are our two variables ? I'm a bit confused

OpenStudy (anonymous):

so you are solving for the cost per miles and for the days

OpenStudy (anonymous):

so m=miles and d=days

OpenStudy (anonymous):

if 300miles per 3 days costs 111 dollars, you can write that algebraically like i did above

OpenStudy (anonymous):

okay and to isolate i would have to multiply both sides by 3m and 5d?

OpenStudy (anonymous):

so first you take one equation, and isolate one variable

OpenStudy (anonymous):

300m/3d=111, becomes 100m/d=111 then multiply by d on both sides, 100m=111d, and divide by 111 on both sides. d=100/111

OpenStudy (anonymous):

now plug that value into the other equation to find the value of m

OpenStudy (anonymous):

okay thank you! that helps me more!